OpenStudy (anonymous):

use the substitution method to solve the following sys. of equations. x+2y-z=2 3x+y-2z=-15 3x+4y-3z=-2

OpenStudy (anonymous):

x = -2y +2 +z y = -3x +2z -15 x = -2(-3x+2x-15) + 2 +z x = 6x -4x +30 + 2 +z x = -32 -z y = -3(-32-z) + 2z -15 y = 96+3z + 2z -15 = 5z + 81 3(-32-z) + 4(5z+81) -3z = -2 Not sure if there are any mistakes

OpenStudy (anonymous):

this is why they invented technology. and matrices

OpenStudy (across):

Linear Algebra says x = -13, y = 2 and z = -11. Use as corroboration.

OpenStudy (anonymous):

what is a corroboration? lol can u tell im horrible at math?

myininaya (myininaya):

\[x+2y-z=2 => x+2y-2=z\] \[3x+y-2z=-15 => 3x+y-2(x+2y-2)=-15\] \[3x+y-2(x+2y-2)=-15=> x-3y+4=-15 => x-3y=-19\] \[x-3y=-19 => x=-19+3y\] \[z=x+2y- 2 \& x=-19+3y => z=-19+3y+2y-2=>z=-19+5y-2\] \[z=-19+5y-2=>z=-21+5y\] \[z=-21+5y \& x=-19+3y \& 3x+4y-3z=-2 =>\] \[3(-19+3y)+4y-3(-21+5y)=-2\] now lets simplify and solve for y \[-51+9y+4y+63-15y=-2\] \[-2y+12=-2 => -2y=-14 => y=7\] i can't figure out where i made mistake i have y=7

myininaya (myininaya):

you should tell your teacher solve this using substitution style is crazy talk lol

OpenStudy (anonymous):

i can solve it by typing http://www.wolframalpha.com/input/?i=x%2B2y-z%3D2+and+3x%2By-2z%3D-15+and+3x%2B4y-3z%3D-2

myininaya (myininaya):

ok i see a mistake i don't know how to multiply

myininaya (myininaya):

3(-19) =-57 not freaking -51

myininaya (myininaya):

so those last two lines should be \[-57+9y+4y+63-15y=-2\] \[-2y+6=-2 => -2y=-8 => y=4\]

OpenStudy (anonymous):

OMG thank you satellite73 and myininaya

myininaya (myininaya):

\[\text{ if y=4 and }z=-21+5y=-21+5(4)=-21+20=-1\] \[\text{ if y=4 and } x=-19+3y=-19+3(4)=-19+12=-7\]

myininaya (myininaya):

yay! :)